Statistics students in Oxnard College sampled 9 textbooks in the Condor bookstore, and recorded number of pages in each textbook and its cost. The bivariate data is shown below,

Number of Pages (xx) Cost(yy)
319 22.95
627 28.35
951 58.55
969 58.45
958 45.9
777 48.85
845 39.25
463 32.15
929 63.45



A student calculates a linear model
y = ? x + ?(Please show your answers to two decimal places)

Use the model above to estimate the cost when number of pages is 742
Cost = $ ?(Please show your answer to 2 decimal places.)
Answer #1

using minitba>stat>regression

we have

Regression Analysis: Cost(y) versus Number of Pages (x)

The regression equation is
Cost(y) = 4.165 + 0.05271 Number of Pages (x)


S = 7.65518 R-Sq = 75.5% R-Sq(adj) = 72.0%


Analysis of Variance

Source DF SS MS F P
Regression 1 1266.34 1266.34 21.61 0.002
Error 7 410.21 58.60
Total 8 1676.55

Ans 1 ) the regression line is

y= 4.17 + 0.05 x

Ans 2 ) when number of pages is 742

Cost(y) =4.165 + 0.05271*742 = 43.28